High / low and least square method

Planters limited produces peanuts. You are the accountant in the production department and have been provided with the following information for the month of June 2018:

Production units                                                               Total Costs $

500                                                                                         3520

440                                                                                         3280

400                                                                                         3120

280                                                                                         3640

160                                                                                         2160

Required:

Use the High Low and Least Squared methods of separating costs to do the following:

(a)   Calculate the variable cost per unit.

(b)   Find the amount of fixed cost present in the cost structure.

(c)   Develop a cost function/equation that relates total cost to activity levels.

(d)   Calculate the production level at which total costs is $2,400.

(e) Calculate the total cost if production level is 1500 units.
Answer #1
Production units Total costs
High level of activity 500 3520
(-) Low level of activity 160 2160
Difference 340 1360
a.
Variable cost per unit = Difference in total costs / Difference in production units = 1360 / 340 4 per unit
b.
Now let us calculate total fixed cost using the data of high level of activity
Total cost = Fixed cost + ( Variable cost per unit * Production in units )
3520 = Fixed cost + ( 4 * 500 )
3520 = Fixed cost + 2000
Fixed cost = 3520 - 2000 1520
c.
Cost function :
y = 1520 + 4x
d.
Here, we are required to calculate 'x' i.e. production units
y = 1520 + 4x
2400 = 1520 + 4x
4x = 2400 - 1520
4x = 880
x = 880 / 4 220 units
e.
Here, we are required to calculate 'y' i.e. total cost
y = 1520 + 4x
y = 1520 + 4*1500 7520